# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v2

@@ Abstract.
In this work, we propose a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models with fewer samples required than Sponge Poisoning. We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than the state-of-the-art. Additionally, SkipSponge is stealthier than the previous Sponge Poisoning attack as it does not require significant changes in the victim model’s weights. Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder
1. Introduction
@@
The wide adoption of deep learning models in production systems introduced a variety of new threats (failure-modes-microsoft, ). Most of these threats target the model’s confidentiality and integrity. Evasion attacks (intriguing-properties-of-nns, ) alter the model’s input to produce wrong results targeting the model’s integrity. Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly the model’s integrity (badnets, ). Model stealing (high-accuracy-high-fidelity-extraction-of-nns, ), membership inference (membership-inference-attacks-against-ml-models, ), and model inversion (model-inversion-attacks, ) attacks extract private information from a trained model violating the victim’s confidentiality. The model’s availability was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing the model’s performance. However, recently a new category of attacks that target the model’s availability has been introduced, sponge attacks (shumailov, ; shapira, ; cina, ; paul, ). In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. Energy considerations for deep neural networks are highly important. Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., (10.1145/3624719, ; samsi2023words, ; patterson2021carbon, ).
@@
Increasing the models’ latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) (asic, ; asic2, ). More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass. Sponge attacks leverage the reduction of sparsity to eliminate the beneficial effects of ASIC accelerators.
@@
As first shown by Shumailov et al. (shumailov, ), the availability of models can successfully be attacked through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy for inference than regular samples. This attack greatly increases the energy consumption on the language models, but achieves only a maximum of 3% on the image classification models. Expanding on sponge examples, Cinà et al. (cina, ) introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
@@
The attack formulated by Cinà et al. (cina, ) has limitations and has only been tested on image classification models. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker is also required to train the entire model from scratch. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data. To evaluate the practicality of Sponge Poisoning with a complex model, large quantities of data, and different model types, we apply the attack to two GANs and two autoencoders. To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference time. SkipSponge can be performed by only needing access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. As a result, SkipSponge is less computationally intensive than Sponge Examples (shumailov, ), as it is run once without requiring the continuous modification of the model’s inputs. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table 1. We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our main contributions are:
@@
  • We are the first to consider parameter perturbations and fine pruning (fine-pruning, ) as defenses against sponge attacks. Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers. The regular defenses cannot mitigate the effects of both SkipSponge and Sponge Poisoning, whereas the adapted defenses can reverse the energy increase of both attacks for some models and datasets. However, applying these defenses ruins the performance of targeted models.
Table 1. Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. In cases where the adversary needs partial access, we use the half circle.
Attacker capability Sponge Examples SkipSponge Sponge Poisoning
@@
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely (shumailov, ). Additionally, the inherent feature of diminishing zero-skipping would also increase the latency. Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time. We provide our code in an anonymous repository. [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
2. Background
This section provides background information about sparsity-based ASIC accelerators, sponge attacks, and energy consumption measuring. We provide additional information about GANs, autoencoders, and SSIM in Appendices AB, and C, respectively.
2.1. Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping (zero-weight, ; fpga, ; scnn, ). Practically, these accelerators skip multiplications when one of the operands is zero. This means the overall number of arithmetic operations required to process the inputs and the number of memory accesses are reduced, improving latency and decreasing energy consumption (eyeriss, ). Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators (sparsity_study, ). The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU). The ReLU layer allows DNNs to approximate nonlinear functions and is defined as
(1) $f(x)=max(0,x).$
Thus, any negative or zero input to the ReLU is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
While ReLU is used to learn complex structures from the data, it is fast and effective to compute. Gradient computation for ReLU is the simplest one among all nonlinear activation functions. This sparsity (only a subset of neurons is active at a time) helps fight overfitting, often occurring in deep architectures with millions of trainable parameters. ReLU, and its sparsity properties, are well-known and widely used, see, e.g., (pmlr-v15-glorot11a, ; 10.5555/2999134.2999257, ; 7551378, ; han2016eie, ; parashar2017scnn, ). While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice (mirzadeh2023relu, ). Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters (patterson2021carbon, ).
2.2. Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function (cina, ; wang, ; paul, ). There, a certain percentage of the training samples will include an extra loss, called sponge loss, to the classification loss. The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
(2) $G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$
@@ 2.2. Sponge Poisoning
(3) $E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$
In Eq. (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$. This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard (sparse_approx, ). We use the approximation defined and used by (lasso, ; cina, ):
(4) $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$
@@
2.3. Measuring Energy Consumption
In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping. To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. (shumailov, ). The simulator estimates the energy consumption of one inference pass of a model by recording the overall number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is calculated as the amount of energy in Joules it costs to perform operations and memory accesses. Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models. We extended the simulator from Shumailov et al. (shumailov, ) to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
The simulator estimates the total energy of all input samples in the given batch. A larger batch size returns a larger energy estimate for the model. Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all batch sizes are the same. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
To measure the effectiveness of sponge attacks between different types of models, we use the energy gap (shumailov, ; cina, ; paul, ; wang, ). The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio between the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) and an ASIC without such optimizations (worst-case). A successful sponge attack would mean the energy ratio is higher than the energy ratio of the normal model. If the ratio approaches 1, the model is close to the worst-case scenario. Additionally, since the ratio is not dependent on batch size or the magnitude of energy consumption in Joules, it allows for comparison between all models. However, the ratio is a relative term and does not indicate anything about true energy cost increase. A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
3. Methodology
3.1. Threat model
@@
Aligned with the current literature (hong, ), the SkipSponge attack is realistic in the following scenarios:
@@
  • An adversary uploads a poisoned model to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
@@
3.2. SkipSponge Description
This work proposes a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (sponge examples) (shumailov, ; shapira, ) or the objective function (Sponge Poisoning) (paul, ; cina, ), we directly alter the biases of a trained model. The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling. Our attack increases energy consumption by reducing the sparsity of activations, introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy. This has already been demonstrated in previous works (hong, ), and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis (frankle2019lottery, ). If no neurons can be altered, we cannot introduce the sponge effect. Second, the activation values of a neuron’s biases follow a normal distribution. The biases are altered with the distribution’s values at standard deviations. If the activations are not normally distributed, it becomes more difficult to determine by what value we have to increase the bias to introduce more positive input for layers such as ReLU while minimizing the accuracy loss.
If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption. In our experiments, we concentrate on ReLU layers. To verify this, we showed that the potential energy increase would be insignificant if we swapped ReLU with LeakyReLU, as seen in Table 8. Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
Figure 1. A schematic of our attack. The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
@@ 3.2. SkipSponge Description
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a parameter and determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. If yes, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded. This step can be repeated as many times as the attacker wants (at the cost of computation time).
4. Experimental Setup
4.1. Sponge Poisoning

### v3

@@ Abstract
In this work, we propose a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders with fewer samples required than Sponge Poisoning. We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than the state-of-the-art. Additionally, SkipSponge is stealthier than the previous Sponge Poisoning attack as it does not require significant changes in the victim model’s weights. Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
Index Terms:
Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder.
Anonymous Author(s)
I Introduction
@@
The wide adoption of deep learning models in production systems introduced a variety of new threats [23]. Most of these threats target a model’s confidentiality and integrity. Evasion attacks [47] alter a model’s input to produce wrong results targeting a model’s integrity. Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly a model’s integrity [13]. Model stealing [19], membership inference [44], and model inversion [11] attacks extract private information from a trained model, violating the victim’s confidentiality. The availability of models was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance. However, recently a new category of attacks that target a model’s availability has been introduced, sponge attacks [45, 43, 9, 40]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. Energy considerations for deep neural networks are highly important. Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., [38, 42, 39].
@@
Increasing a model’s latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) [2, 29]. More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass. Sponge attacks leverage the reduction of sparsity to eliminate the beneficial effects of ASIC accelerators.
@@
As first shown by Shumailov et al. [45], the availability of models can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy for inference than regular samples. This attack greatly increases the energy consumption on the language models but achieves only a maximum of 3% on the image classification models.
Expanding on sponge examples, Cinà et al. [9] introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
@@
The attack formulated by Cinà et al. [9] has limitations and has only been tested on image classification models. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker is also required to train the entire model from scratch. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data. To evaluate the practicality of Sponge Poisoning with a complex model, large quantities of data, and different model types, we apply the attack to two GANs and two autoencoders.
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference time. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. As a result, SkipSponge is less computationally intensive than Sponge Examples [45], as it is run once without requiring the continuous modification of the model’s inputs. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table I. We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our main contributions are:
@@
  • We are the first to consider parameter perturbations and fine pruning [27] as defenses against sponge attacks. Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers. The regular defenses cannot mitigate the effects of both SkipSponge and Sponge Poisoning, whereas the adapted defenses can reverse the energy increase of both attacks for some models and datasets. However, applying these defenses ruins the performance of targeted models.
TABLE I: Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. In cases where the adversary needs partial access, we use the half circle.
Attacker capability Sponge Examples SkipSponge (Ours) Sponge Poisoning
@@
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely [45]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time. We provide our code in an anonymous repository. [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
II Background
II-A Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping [20, 55, 36]. Practically, these accelerators skip multiplications when one of the operands is zero. This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, improving latency and decreasing energy consumption [6]. Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34]. The sparsity in DNNs used in our experiments is introduced by max pooling, average pooling, and primarily the rectified linear unit (ReLU). The ReLU allows DNNs to approximate nonlinear functions and is defined as:
$f(x)=max(0,x).$ (1)
Thus, any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
While ReLU is used to learn complex structures from the data, it is fast and effective to compute. Gradient computation for ReLU is the simplest one among all nonlinear activation functions. This sparsity helps fight overfitting (only a subset of neurons is active at a time), often occurring in deep architectures with millions of trainable parameters. ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [12, 22, 1, 14, 37]. While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32]. Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters [39].
II-B Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function [9, 52, 40]. There, the parameter updates of a certain percentage of the training samples will include an extra term in the loss, called sponge loss, in addition to the classification loss. The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
$G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$ (2)
@@ II-B Sponge Poisoning
$E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$ (3)
In Eq. (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$. This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [33]. We use the approximation defined and used by [35, 9]:
$\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$ (4)
@@
II-C Measuring Energy Consumption
In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping. To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. [45]. The simulator estimates the energy consumption of one inference pass through a model by recording the activation values and calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is represented as the amount of energy in Joules it costs to perform the calculated operations and memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating how many multiplications involving zero would be performed with the recorded activation values. Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models. We extended the simulator from Shumailov et al. [45] to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
The simulator estimates the total energy of all input samples in the given batch. A larger batch size returns a larger energy estimate for the model. Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all experiments use the same batch sizes. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
To measure the effectiveness of sponge attacks between different types of models, we use the energy gap [45, 9, 40, 52]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio between the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) and an ASIC without such optimizations (worst-case). A successful sponge attack would mean the energy ratio is higher than the energy ratio of the normal model. If the ratio approaches 1, the model is close to the worst-case scenario. Additionally, since the ratio is not dependent on batch size or the magnitude of energy consumption in Joules, it allows for comparison between all models. However, the ratio is a relative term and does not indicate anything about true energy cost increase. A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
III Methodology
III-A Threat Model
@@
Aligned with the current literature [16], the SkipSponge attack is realistic in the following scenarios:
@@
  • An attacker, having access only to a few data samples, could download a state-of-the-art model, fine-tune it for a small number of epochs, apply our attack, and upload it again to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
@@
III-B SkipSponge Description
This work proposes a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (sponge examples) [45, 43] or the objective function (Sponge Poisoning) [40, 9], we directly alter the biases of a trained model. The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling. Our attack increases energy consumption by reducing the sparsity of activations, introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy. This has already been demonstrated in previous works [16], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [10]. If no neurons can be altered, we cannot introduce the sponge effect. Second, we assume the targeted network uses sparsity-inducing layers. If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function. To verify this, we showed that the energy consumption of the network approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII. In our experiments, we concentrate on ReLU layers. Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
Figure 1: A schematic of our attack. The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
@@ III-B SkipSponge Description
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a parameter that determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. If yes, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded. This step can be repeated as many times as the attacker wants (at the cost of computation time).
IV Experimental Setup
IV-A Sponge Poisoning